A cabin cruiser leaves a harbor and travels to a small island at an average speed of 24 mph. On the return trip, the cabin cruiser travels at an average speed of 16 mph. The total time for the trip is 4 hours. How far is the island from the harbor?

Answer by richwmiller(17219)

x+y=4,
24x=16y
x = 8/5= 1.6, y = 12/5 =2.4
distance = 192/5=38.4